# Token Spammers, Rug Pulls, and SniperBots: An Analysis of the Ecosystem of Tokens in Ethereum and the Binance Smart Chain (BNB)

### v2

@@ 11 Ethical considerations
In this work, we analyzed two blockchains, Ethereum and BNB smart chain, and processed over 3 billion transactions. We looked at the addresses that created tokens on the two blockchains and attempted to understand how the tokens are used. By design, all the data are publicly stored in the blockchain ledger, and the EOAs addresses are pseudo-anonymous. During our study, we never attempt to correlate users’ addresses with external events jeopardizing the privacy of the users. Moreover, we never collected nor attempted to retrieve the IP addresses of the users. Consequently, according to our IRB’s policy, we did not need any explicit authorization to perform our experiments.
@@ 12 Conclusion and Future Work
In this work, we thoroughly investigate the tokens and the liquidity pools of the BNB Smart Chain and Ethereum. Although the BSC is five years younger than Ethereum, we find in BSC three times more tokens and liquidity pools than in Ethereum. One of the most surprising differences between the two blockchains is the remarkable difference in the purpose of deployed smart contracts. Indeed, in BSC 61% of smart contracts are token contracts, while in Ethereum, only the 8.5%.
We studied the lifetime of the tokens and who generates them. Here, we discovered two very interesting metrics: 60% of the total tokens of both blockchains do not survive to their first day (1-day token), and a tiny fraction of addresses (1% of addresses), which we called token spammers, created more than 20% of the tokens. We explore the correlation between token spammers and 1-day tokens, and we found that token spammers strongly impact the existence of 1-day tokens. More interestingly, we find that token spammers use 1-day tokens as disposable tokens to arrange frauds exploiting the mechanism of liquidity pools.
We selected from our dataset all the liquidity pools that show evidence of the frauds and used the retrieved data to dissect them, analyzing them from several perspectives. Finally, we introduce the sniper bot, trading bots that aim to buy tokens at their listing price. However, they unwillingly became victims of the scam because of their mechanism.
As future work, we believe it is interesting to further refine our results by verifying and including addresses that cooperate to perpetrate exit scams in the analysis. Moreover, it could be possible to search also for other malicious and more sophisticated patterns. Finally, another promising direction is further exploring sniper bots to provide a more detailed analysis of their typologies and operations.
Appendix A Appendix

### v3

@@ 12 Ethical considerations
In this paper, we examined 3 billion transactions from Ethereum and the BSC. We focused our research on the addresses that create tokens and how they use them. All data we retrieved is publicly available, and EOA addresses are pseudo-anonymous. We never attempted to deanonymize the addresses or violate their privacy during this work. Consequently, and in accordance with our IRB’s policies, we did not require express approval to conduct our analysis.
@@ 13 Conclusion and Future Work
In this work, we conduct a thorough investigation of the tokens and the liquidity pools of the BNB Smart Chain and Ethereum. We studied the lifetime of the tokens and their creators. We discovered two very interesting metrics: 60% of the total tokens of both blockchains do not survive their first day (1-day token), and a tiny fraction of addresses (1% of addresses), which we called token spammers, created more than 20% of the tokens. We explore the correlation between token spammers and 1-day tokens, and we found that token spammers strongly impact the existence of 1-day tokens.
More interestingly, we find that token spammers use 1-day tokens as disposable tokens to arrange rug pulls, exploiting the mechanism of liquidity pools. We selected from our dataset all the liquidity pools that show evidence of a rug pull and dissect the operations, analyzing them from several perspectives. Finally, we introduce the sniper bot, trading bot that aims to buy tokens at their listing price. However, they unwillingly became victims of the rug pulls because of their mechanism.
As future work, we believe it is interesting to further refine our results by including addresses that cooperate to perpetrate rug pulls in the analysis. It could be possible to uncover other malicious and more sophisticated patterns. As discussed in Sec. 11, cost-efficient blockchains could be more exposed to the 1-day rug pulls. Thus, it is interesting to extend our analysis to blockchains with transaction costs comparable to BSC (e.g., Algorand [30]). Finally, another promising direction is further exploring sniper bots to provide a more detailed analysis of their typologies and operations.
Acknowledgments
This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. This work was partially supported by the MIUR under grant “Dipartimenti di eccellenza 2018-2022" of the Department of Computer Science of Sapienza University.
Appendix A Functions and events of the ERC-20 and BEP-20 standards
@@
Appendix B Are 1-day rug pulls frauds?
1-day rug pulls are very different from more notorious rug pulls like Squid Game [47] or Luna Yield [17]. Indeed, these operations lasted weeks or months, and their perpetrator exploited extensive marketing campaigns and misleading advertising to deceive users into investing in their tokens. In the case of Squid Game, the scammer created a token in the BSC following the success of the homonym Netflix television series [54]. Due to the extensive marketing campaign promoting the token as official on social media platforms such as Twitter and Telegram, its value skyrocketed from a few cents to over \$2,856 in less than a week [19]. Then, the scammer removed nearly all of the liquidity from the pool (\$3.3 million), causing the token’s value to plummet to near zero [47]. In our paper, we study 1-day operations that aim to make a profit with the least possible effort in a short time frame. For this reason, it is unlikely that they leverage sophisticated marketing campaigns to lure investors, like in the case of Squid Game. However, some 1-day rug pull operations use other kinds of deceptive tactics. The first uses token names identical or slightly different from well-known companies or popular tokens. As we saw in Sec. 7.2.3, this case involves 8.7% of Ethereum rug pulls, and 10% of BSC rug pulls. Another deceptive technique consists in attempting to legitimate the project by verifying the smart contract code on BSCscan and Etherscan. The verification consists in uploading the source code so that the platform can compile it and verify that it matches the bytecode of the token stored in the blockchain. The verification provides users transparency and gives more guarantee that the token is not fraudulent. We find the smart contract is available and verified for 55% (147,069) of BSC and 67% (14,722) of Ethereum tokens involved in the 1-day rug pull operations. Finally, another technique to legitimate the token consists in creating the "official" Telegram group of the token. We find evidence of this technique in the smart contract’s code and then inspect the groups on Telegram. Indeed, analyzing the source codes, we notice that 19,096 token smart contracts in BSC and 1,334 in Ethereum report a link to the Telegram group of the token. Although the organizers of 1-day rug pulls use deceptive techniques to dupe investors, we cannot consider these operations frauds because the phenomenon is still not regulated. In any case, people lose money: Investors bought the token in 92.7% of the Ethereum rug pulls and in 91.2% of the BSC ones, and in all these cases the investment is lost. For this reason, we believe that these operations, even if not illegal, are exploitative of the DeFi ecosystem and should be contrasted to safeguard investors. Indeed, regulators are starting to take action to contrast them. For example, New York State Senator Kevin Thomas proposes criminalizing rug pulls and other crypto frauds by introducing a new bill amendment request (Senate Bill S8839) [55]. The idea of the bill is to introduce the crime of illegal rug pull that occurs if the creator of the token sells more than 10% of his tokens within five years of their last sale.
Appendix C Token names
Table 6: Token names most frequently used in 1-day rug pull operations.
BNB Smart Chain Ethereum
Name # of tokens Name # of tokens
Pornhub 1,023 Hyve.works 50
Galaxy 588 Deriswap 32
Seedswap 502 Shibaswap 28
Lionswap 429 Apple core finance 17
Eco.finance 421 X20.finance 16
Spacex 419 Yield farm rice 15
Onlyfans 419 The sandbox 14